1. Oxygen Supply in Submarines Nuclear submarines can stay under water nearly indefinitely because they can produce their own oxygen by the electrolysis of water.
How many liters of O2 at 298 K and 1.00 bar are produced in 2.25 hr in an electrolytic cell operating at a current of 0.0400 A? answer in L
1
Expert's answer
2020-05-01T14:41:03-0400

m=8 x 2.25 x 0.04 / 26.8 = 0,0268656716 g.

n=m/M = 0,0268656716/32 = 0,000839552238.

pV= nRT. V= nRT/p = 0,000839552238 x 8.314 x 298 / 100 = 0,0208005112 L.